OpenStudy (anonymous):

Translate the sentence into an equation. Four times a number is 24. A. 4n = 24 B. 4 + n = 24 C. D. n = 4 ∙ 24

OpenStudy (sleepyjess):

Welcome to OpenStudy! What do you think?

OpenStudy (anonymous):

i like it i think that it helps and teaches people to work together.

OpenStudy (sleepyjess):

huh? I meant what do you think about your question?

OpenStudy (sleepyjess):

But yes, you are correct about that OpenStudy helps and teaches people to work together ^_^

OpenStudy (anonymous):

oh sorry. i don't know am not really good with math.

OpenStudy (sleepyjess):

Ok, that is fine :) 4 times a number is 4*n, or 4n. is 24, this mean equals 24.

OpenStudy (anonymous):

i think it is A

OpenStudy (wolfgirl12):

It's A. 4n=24.

OpenStudy (anonymous):

i have 5 more qestions

OpenStudy (wolfgirl12):

okay what are they?

OpenStudy (sleepyjess):

Okie ^_^

OpenStudy (anonymous):

Translate the sentence into an equation. The product of 8 and a number increased by 4 is 60. A. 8y + 4 = 60 B. 4y + 8 = 56 C. 8 ÷ y + 4 = 56 D. y + 8 + 4 = 56

OpenStudy (sleepyjess):

What do you think?

OpenStudy (wolfgirl12):

It's A. 8y+4=60 A number (y) times 8 and then +4 =60

OpenStudy (sleepyjess):

@wolfgirl12 , Please do not give out direct answers.

OpenStudy (anonymous):

i think it is A because its the only one that equals 60

OpenStudy (sleepyjess):

Yes, but that isn't a very good reason ^_^ the product of 8 and a number, 8y, increased by 4, + 4, is 60, = 60.

OpenStudy (anonymous):

oh okay

OpenStudy (sleepyjess):

Could you put what you think on each when you post them?

OpenStudy (anonymous):

Translate the sentence into an equation. Seven times the difference of a number and 3 is 35. A. 7k – 3 = 35 B. k – 3 ∙ 7 = 35 C. 7 ∙ (k – 3) = 35 D. 7 ∙ (k + 3) = 35

OpenStudy (anonymous):

does it look okay on you guys screen

OpenStudy (sleepyjess):

Just refresh

OpenStudy (sleepyjess):

What do you think on this one?

OpenStudy (anonymous):

i was going to say it is c but i really think it is a what do you think

OpenStudy (sleepyjess):

Sorry, trying to help multiple people at once haha

OpenStudy (sleepyjess):

7 times, 7*, the difference of a number and 3, (k - 3), is 35, = 35

OpenStudy (anonymous):

Translate the sentence into an equation. Four times the difference of a number and 5 is 24. A. 4 • (c – 5) = 24 B. 4c – 5 = 24 C. c – 4 • 5 = 24 D. 5 • (c – 4) = 24

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

what is the answer to this one